Calculate the specific volume of normal butane at a pressure of 80 bar and a temperature of 227° C is desirable as follows:



1, the ideal gas equation



2, the compressibility coefficient of the Peters method



3, virial equations up to a factor of two



4, Redlish Quang equation

1
Expert's answer
2022-01-23T17:36:02-0500

1. The ideal gas equation:


"pV=\dfrac mMRT→v=\dfrac Vm=\dfrac{RT}{pM},\\\space\\
v=8.84·10^{-3}\text{ m}^2/\text{kg}."


2. The compressibility factor:


"z=\dfrac{pV}{nRT}=0.99973."


The specific volume:


"v_P=zv=8.838·10^{-3}\text{ m}^2/\text{kg}."


3. Virial equations up to a factor of two:


"Z\equiv\dfrac{Pv}{RT}=17.23."


4. The coefficients for Redlish-Quang equation (the critical pressure for butane is 3.796 MPa, critical temperature for butane is 425.1 K):


"a=0.42748\dfrac{R^2T_c^{2.5}}{P_c}=29.0059,\\\space\\
b=0.08664\dfrac{RT_c}{P_c}=29.0059=8.067·10^{-5}."